4.19 If P(A) 0.7, P(B)0.6, and A and B are independent, find P(A and B). PREFE 4.20 If P(A) 0.3, P(B)0.4, and P(A and B) 0.2, are A and B independent? Name-E Store B
Answer #1

9) If A and B are independent events, P(A and B) = P(A) * P(B) = 0.7 * 0.6 = 0.42

10) P(A) * P(B) = 0.3 * 0.4 = 0.12

Since, P(A and B) is not equal to P(A) * P(B) , A and B are not independent events.